1. Find the slope of the line containing the points (3,1) and (-4, -1).
2. Determine the equation of the line that passes through the points (-4, -3) and (2,2)

Answer by checkley71(8403)

THE SLOPE THROUGH THESE TWO POINTS IS (-1-1)/-4-3) OR -2/-7 OR 2/7 FOR THE SLOPE.
NOW SUBSTITUTING ONE OF THE SET OF POINTS INTO THE EQUATION Y=mX+b WE GET 1=(2/7)*3+b OR 1=6/7+b OR b=1-6/7 OR b=1/7. THUS THE EQUATION OF THIS LINE IS Y=6/7X+1/7